OpenStudy (eratraya):

Which of the following is a rational number? square root 15 square root 16, square root 21, square root 22 square root 15 square root 16 square root 21 square root 22

OpenStudy (eratraya):

@zepdrix do you think you can help

OpenStudy (loser66):

What is a rational number?

OpenStudy (eratraya):

a rational number is a number that can be written as a fraction a/b or a repeating decimal or terminating decimal

OpenStudy (loser66):

ok, so \(16 = \color{red}{?}^2\)

OpenStudy (eratraya):

8?

OpenStudy (loser66):

\(8^2=8*8 =64\neq 16\)

OpenStudy (eratraya):

so its b

OpenStudy (loser66):

ok!! if you just want to guess, it's ok to me. And your guessing is correct.

OpenStudy (eratraya):

sweet can you help me with a couple more

OpenStudy (eratraya):

@Loser66